Rates for a couple filing a joint tax return are 10% up to $17,850 and 15% thereafter up to $72,500. Joanna and Delbert Linton earned combined gross income of $26,000. They took the standard deduction of $12,200 for a married couple filing a joint return and two exemptions of $3,900 each.

Compute the amount of income tax.
Answer #1

Answer:

Filing status: MFJ

Gross income =  $26,000

Less, Standard Deduction = $12,200

Less, Personal exemptions = $3,900 * 2 = $7,800

Taxable Income = 26000 - 12200 - 7800 =$6000

Rates for a couple filing a joint tax return are 10% up to $17,850 and 15% thereafter up to $72,500.

Tax liability = 6000 * 10 = $600

Amount of income tax = $600